### v6

# CompMarkGS: Robust Watermarking for Compressed 3D Gaussian Splatting

Abstract
As 3D Gaussian Splatting (3DGS) is increasingly adopted in various academic and commercial applications due to its high-quality and real-time rendering capabilities, the need for copyright protection is growing. At the same time, its large model size requires efficient compression for storage and transmission. However, compression techniques, especially quantization-based methods, degrade the integrity of existing 3DGS watermarking methods, thus creating the need for a novel methodology that is robust against compression. To ensure reliable watermark detection under compression, we propose a compression-tolerant 3DGS watermarking method that preserves watermark integrity and rendering quality. Our approach utilizes an anchor-based 3DGS, embedding the watermark into anchor attributes, particularly the anchor feature, to enhance security and rendering quality. We also propose a quantization distortion layer that injects quantization noise during training, preserving the watermark after quantization-based compression. Moreover, we employ a frequency-aware anchor growing strategy that enhances rendering quality by effectively identifying Gaussians in high-frequency regions, and an HSV loss to mitigate color artifacts for further rendering quality improvement. Extensive experiments demonstrate that our proposed method preserves the watermark even under compression and maintains high rendering quality.
1 Introduction
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has recently emerged as an impactful method for novel view synthesis, offering real-time and photorealistic rendering capabilities. Since its introduction, 3DGS has undergone rapid advancements in the field of image-based 3D reconstruction (Charatan et al., 2024; Lin et al., 2024; Zhang et al., 2024), attracting significant attention from academia and industry. One drawback of 3DGS is that it is composed of a vast number of 3D Gaussians, requiring substantial storage space. This limitation has motivated the development of 3DGS compression techniques to reduce storage requirements and facilitate efficient transmission of 3DGS models. In parallel, its growing commercial adoption in areas such as digital twin construction and AR/VR applications has raised critical concerns regarding the copyright protection of trained 3DGS models. These concerns have led to increased research interest in 3DGS watermarking. However, to the best of our knowledge, no prior work has specifically addressed the challenge of creating a 3DGS watermark that can withstand model compression, a scenario illustrated in Fig. 1.
Existing 3DGS watermarking methods (Chen et al., 2024b; Guo et al., 2024; Huang et al., 2025; Jang et al., 2025; Tan et al., 2024) embed watermarks by directly modifying Gaussian attributes, which makes the watermark vulnerable to loss when those attributes are altered during compression. In particular, quantization‐based compression shifts the distribution of model parameters during the compression process, severely degrading watermark performance. As a result, existing methods fail to provide adequate copyright protection for compressed 3DGS models.
Figure 1: Application scenario of CompMarkGS. Owners embed the watermark into a 3DGS asset using our proposed method, then compress and distribute the model across various digital environments. The owner can extract the hidden message to verify ownership even if an unauthorized user alters the distributed model or its rendered images.
To handle this problem, we propose CompMarkGS, an anchor-based 3DGS watermarking method. Our method leverages the anchor-based 3DGS architecture, which dynamically predicts Gaussian attributes from anchor points via multiple implicit multi-layer perceptrons (MLPs). This implicit representation makes it extremely difficult for a malicious user to detect the presence of a watermark by directly analyzing the point cloud or Gaussian attributes. We seamlessly integrate the watermark by embedding a learnable watermark feature into the anchor feature, which is one of the anchor attributes, without altering the original model architecture. Unlike prior work focusing on image-domain distortions, we tackle the novel challenge of watermark robustness against model quantization. We propose a quantization distortion layer (QDL) that injects quantization noise into anchor attributes during training, teaching the watermark to be resilient. As a result, our model achieves both high rendering quality and robust watermark integrity after compression.
To further improve rendering quality, we introduce a frequency-aware anchor growing strategy. Specifically, we identify high-frequency regions in the rendered image, map them to Gaussian coordinates, and selectively apply anchor growing in those regions. In addition to this structural enhancement, we design an HSV loss to address color artifacts that occur when embedding the watermark. This loss function operates by constructing a binary mask over regions with noticeable color artifacts, identified by analyzing the hue (H), saturation (S), and value (V) components. This mask guides the loss computation by focusing on these specific regions, improving the rendering quality.
Our extensive experiments on the Blender (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022) datasets show that CompMarkGS preserves watermark integrity and rendering quality more effectively both before and after compression. Each component of our method—anchor-based 3DGS watermarking, the quantization distortion layer, and the frequency-aware anchor growing strategy—independently enhances watermark fidelity while preserving rendering quality, even under compression and various watermarking attacks. Furthermore, our method outperforms state-of-the-art approaches across various watermark message lengths. Our main contributions are as follows:
  • We propose CompMarkGS, a compression-tolerant anchor-based watermarking method for 3DGS that embeds a learnable watermark embedding feature into the anchor feature, preserving the model structure while ensuring high rendering quality and bit accuracy.
  • We introduce a quantization distortion layer that injects quantization noise during training, enabling 3DGS watermarking to remain robust against quantization-based compression while maintaining high rendering quality.
  • We propose a frequency-aware anchor growing strategy that selectively expands anchors in high-frequency regions, which enhances rendering quality, and an HSV loss that mitigates color artifacts to further improve rendering quality.
2 Related Works
3D Gaussian Splatting representation.
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) is an innovative 3D representation technique that models scenes using explicit primitives. It renders images by projecting these primitives into arbitrary viewpoints and seamlessly blending pixel colors via an $\alpha$-blending, achieving high-quality results while substantially accelerating real-time rendering. Thanks to these advantages, 3DGS has been widely adopted in various research areas, including avatars (Abdal et al., 2024; Chen et al., 2025a; Qian et al., 2024; Yuan et al., 2024), dynamic scenes (Li et al., 2024; Wu et al., 2024; Yan et al., 2025; Zhu et al., 2025), and 3D generation (Tang et al., 2023; Xie et al., 2024b; Yi et al., 2024; Zhou et al., 2024). Recently, Scaffold-GS (Lu et al., 2024) has advanced the 3DGS framework by introducing anchor points to construct a hierarchical 3D representation. Anchor-based methods (Lee et al., 2024b; Ren et al., 2024; Wang et al., 2024b) effectively minimize redundant Gaussians, enhancing rendering quality and increasing robustness against view changes. This representation approach is promising for 3DGS compression optimization, as it reduces parameter counts while maintaining high-quality rendering.
3D Gaussian Splatting compression.
To achieve high-quality rendering images, 3DGS generates a considerable number of Gaussians, leading to significant storage overhead. To address the storage overhead, vector quantization methods (Fan et al., 2025; Lee et al., 2024a; Navaneet et al., 2024; Xie et al., 2024a) have been widely explored. These methods prune Gaussians with minimal impact on rendering quality and use codebooks to encode the attributes of Gaussians compactly. More recently, anchor-based representations have gained significant attention in 3DGS (Chen et al., 2024a; Lu et al., 2024; Wang et al., 2024a). Compression methods (Chen et al., 2025b; 2024a; Wang et al., 2024a; Zhan et al., 2025) that leverage the structural relationships between anchors have shown superior performance. For example, HAC (Chen et al., 2024a) reduces spatial redundancies among anchors by utilizing hash grids for parameter quantization and enabling entropy modeling. ContextGS (Wang et al., 2024a) introduces a unified compression framework with a factorized prior, enabling entropy modeling of anchor features while leveraging hierarchical anchor relations to reduce redundancy among anchors. CAT-3DGS (Zhan et al., 2025) projects Gaussian primitives onto PCA-aligned triplanes and applies spatial autoregressive coding to capture spatial correlations, thereby enhancing entropy coding efficiency. Given their compression efficiency, anchor-based representations are especially advantageous for real-world applications.
Steganography and digital watermarking.
Information hiding research is primarily divided into two categories: digital watermarking and steganography, which differ in their core objectives. Steganography aims to conceal information in digital assets with invisibility as the primary metric. Driven by advances in deep learning, various methods (Tancik et al., 2020; Biswal et al., 2024) have been proposed to hide information in media such as images and videos, with recent work (Dong et al., 2024; Li et al., 2023) extending into 3D scene representation. For example, GS-Hider (Zhang et al., 2025a) introduces the first steganography technique for 3D Gaussian Splatting (3DGS), and SecureGS (Zhang et al., 2025b) proposes an anchor-based steganographic framework that hides information using a private MLP. In contrast, digital watermarking embeds ownership data to protect assets and prioritizes robustness to distortions. Early watermarking research focused on the pixel domain (Van Schyndel et al., 1994; Wolfgang & Delp, 1996) or the frequency domain (Barni et al., 2001; Navas et al., 2008). Subsequently, deep learning-based methodologies (Luo et al., 2020; Zhu et al., 2018) demonstrated superior robustness against a variety of distortions. This research has recently expanded to protect radiance field models like NeRF (Mildenhall et al., 2021). WateRF (Jang et al., 2024) proposes a plug-and-play method that uses the frequency domain for robust watermarking. With the advent of 3DGS, GaussianMarker (Huang et al., 2025) embeds the watermark by estimating Gaussian uncertainties and integrating specific Gaussians. 3D-GSW (Jang et al., 2025) embeds the watermark by selectively removing Gaussians to minimize the impact on rendering quality. GuardSplat (Chen et al., 2024b) introduces an SH‐aware embedding mechanism. However, these works only consider traditional distortions and do not account for real-world attacks such as model compression. In this paper, we propose a watermarking method that is robust to model compression.
3 Preliminaries
Scaffold-GS.
Scaffold-GS (Lu et al., 2024) clusters adjacent Gaussians using anchor points, reducing redundancy. Initial anchor points are placed at voxel centers, and the attributes for each anchor consist of an anchor feature $f\in\mathbb{R}^{d}$, scaling factor $l\in\mathbb{R}^{3}$, and $K$ learnable offsets ${O}\in\mathbb{R}^{K\times 3}$, where $K$ is the number of Gaussians. Visible anchor points within the viewing frustum generate $K$ Gaussians, with their positions computed as follows:
$\mu_{k}=x_{a}+O_{k}\odot l,$ (1)
where $x_{a}\in\mathbb{R}^{3}$ is the anchor point position, $O_{k}\in\mathbb{R}^{3}$ denotes the $k$-th offset vector, $\mu_{k}$ is the generated $k$-th Gaussian position, and $\odot$ represents the element-wise product. The Gaussian attributes opacity $\alpha_{i}$, color $c_{i}$, quaternion $q_{i}$ and scale $s_{i}$ are predicted using separate MLPs:
$\{\alpha_{i},c_{i},q_{i},s_{i}\}^{K}_{i=1}=\text{MLP}(f,\delta_{av},\vec{\mathbf{d}}_{av}),$ (2)
where $\delta_{av}=\parallel{x}_{a}-{x}_{v}\parallel_{2}$ and $\vec{\mathbf{d}}_{av}=\frac{{x}_{a}-{x}_{v}}{\parallel{x}_{a}-{x}_{v}\parallel_{2}}$ denote the relative distance and viewing direction between the camera ($x_{v}$) and the anchor ($x_{a}$) positions.
Figure 2: Overview of CompMarkGS. A learnable watermark embedding feature $f^{w}$f is added to the anchor feature $f$ to create a watermarked anchor feature $f^{w}$. The quantized watermarked anchor feature from the quantization distortion layer is then fed into individual MLPs to predict Gaussian attributes. The watermark is extracted from the low-frequency band of the rendered image via a pre-trained decoder. A frequency-aware anchor growing strategy identifies Gaussians in high-frequency regions to introduce additional anchors. Moreover, an HSV loss $\mathcal{L}_{hsv}$, and a frequency loss $\mathcal{L}_{freq}$, are applied to mitigate visual artifacts. The entire model is optimized via the total loss $\mathcal{L}_{total}$.
4 Method
4.1 Anchor-based 3DGS watermarking
Anchor-based 3DGS (Lu et al., 2024) enhances the security of embedded watermarks by hiding explicit Gaussian attributes behind anchor features processed by implicit MLPs. This property of anchor-based 3DGS makes it more difficult for attackers to directly manipulate embedded watermarks and Gaussian attributes. Moreover, the anchor feature $f$ influences the rendering quality indirectly through MLPs rather than directly controlling Gaussian attributes. On the other hand, attributes such as position $x$, scaling factor $l$, and offsets $O$ directly determine Gaussian attributes, and modifying them leads to visible distortions in the rendered images. To embed a watermark into the anchor feature while preserving rendering quality, we introduce a learnable watermark embedding feature $f^{\prime}\in\mathbb{R}^{d}$, with the same dimension as the anchor feature $f\in\mathbb{R}^{d}$. This design aims to achieve an optimal balance between rendering quality and watermark robustness. The resulting watermarked anchor feature $f^{w}$ is defined as follows:
$f^{w}=f+\mathrm{tanh}(f^{\prime}),\quad\mathrm{where}\quad f,f^{\prime}\in\mathbb{R}^{d},$ (3)
where the watermarked anchor feature $f^{w}$ is generated by the element-wise sum of the anchor feature $f$ and a learnable watermark embedding feature $f^{\prime}$. The term $\mathrm{tanh}(\cdot)$ denotes the element-wise hyperbolic tangent function, which we apply to the watermark embedding feature $f^{\prime}$ to bound its values within the range [1,1][-1,1] before addition. This bounding process is crucial for stabilizing gradients and preserving rendering quality. In a well-trained anchor-based 3DGS model, the components of the anchor attributes are known to follow a normal distribution (Lu et al., 2024). Adding an unbounded feature would increase the variance of this distribution, leading to unstable gradients, slower convergence, and degraded rendering quality. While a sigmoid function could be used, its $[0,1]$ output range would introduce a directional bias by only adding positive values. The symmetric [1,1][-1,1] range of the tanh function prevents this bias, ensuring the watermark is added with a consistent magnitude. As a result, our method achieves stable convergence and efficient embedding without degrading rendering quality. A detailed analysis is provided in Appendix E.1.
4.2 Robustness to quantization-based compression
In digital watermarking research, differentiable distortion layers have long been used to enhance robustness by simulating traditional distortions, such as JPEG and cropping, during training (Zhu et al., 2018). However, this method has not been applied in existing 3DGS watermarking research to the problem of model compression, leading to significant watermark loss during quantization.
To overcome this limitation, we propose a quantization distortion layer (QDL) that simulates quantization-induced distortions, encompassing losses in watermark robustness and rendering fidelity. Our QDL leverages quantization techniques from prior anchor-based 3DGS compression methods (Chen et al., 2024a). While the original method uses this mechanism for model compression, our work is the first to adapt it as a form of data augmentation to train a robust watermark. During watermark training, the QDL simulates the quantization process by injecting noise into the watermarked anchor feature $f^{w}$. By integrating the QDL into the embedding stage, the watermark is thus optimized to remain robust against subsequent compression attacks. The quantized watermarked anchor feature for the $i$-th anchor $\tilde{f}^{w}_{i}$ is defined as follows:
$\tilde{f}^{w}_{i}=f^{w}_{i}+$ $\mathcal{U}\Bigl(-\frac{1}{2},\frac{1}{2}\Bigr)\cdot q_{i},\quad\text{where}\quad q_{i}=\mathrm{Q_{0}}\cdot\Bigl(1+\tanh(r_{i})\Bigr),\quad r_{i}=\mathrm{MLP}_{q}(f^{w}_{i}),$ (4)
where $q_{i}\in\mathbb{R}$ denotes the quantization scale restricted to $0,2\mathrm{Q_{0}}]$, and $r_{i}\in\mathbb{R}$ is the refinement output from an $\text{MLP}_{q}$ that takes $f^{w}_{i}$ as input to adjust the initial quantization scale $\mathrm{Q_{0}}$. By scaling a randomly generated $d$-dimensional noise vector from a uniform distribution over $\frac{1}{2},\frac{1}{2}]$ by $q_{i}$ and adding it to the watermarked anchor feature $f^{w}$, we simulate the rounding errors caused by quantization. This process forces the watermark to become resilient to the specific distortions it will encounter during a real compression attack.
4.3 Frequency-aware anchor growing
We apply a Discrete Wavelet Transform (DWT) to the rendered image and feed its low-frequency Low-Low ($LL$) subband, $I_{LL}$, into the pre-trained HiDDeN (Zhu et al., 2018) decoder $D$ to extract the watermark message $M^{\prime}=D(I_{LL})$. The low-frequency band is ideal for embedding watermarks to ensure robustness, but it is also critical for overall quality as it contains the image’s global structure. This creates a fundamental conflict. If we attempt to embed a watermark and enhance quality in the same low-frequency band, the quality enhancement algorithm may treat the watermark as unwanted noise and remove it, thereby degrading the watermark’s robustness.
Based on this insight, we propose a frequency-aware anchor growing (FAG) strategy. To selectively grow anchors in detail-rich, high-frequency regions, our method first transforms the rendered image $I^{\prime}$ and the ground truth image $I$ into the frequency domain using the Discrete Fourier Transform (DFT), as shown in Fig. 2. We then apply a high-pass filter to isolate the high-frequency components and reconstruct the corresponding high-frequency images $I^{\prime}_{hf}$, and $I_{hf}$ using the Inverse Discrete Fourier Transform (IDFT). To generate a binary mask that selects pixels in these high-frequency regions, we define the pixel-wise SSIM loss $P_{error}$ as follows:
$P_{error}(p)=1-\mathrm{SSIM}(I_{hf}(p),I^{\prime}_{hf}(p)),$ (5)
where $p$ is a pixel coordinate, $I^{\prime}_{hf}(p)$ is the pixel value in the high-frequency region of the rendered image, and $I_{hf}(p)$ is the corresponding value in the ground truth. After computing the median of all pixel-wise SSIM losses, denoted as $\tilde{P}_{error}$, we create a binary mask by selecting pixels within the range $\tilde{P}_{error}-\epsilon,\tilde{P}_{error}+\epsilon]$. This mask is then used to extract the 2D coordinates of high-frequency regions. Subsequently, each 3D Gaussian is projected onto the 2D image plane, and its continuous coordinates are rounded to obtain integer pixel‐level coordinates. By matching these 2D Gaussian coordinates with the high-frequency pixel coordinates, we generate a boolean mask $F_{mask}$ that identifies anchors located in the high-frequency regions. Only the Gaussians selected by $F_{mask}$ are activated for the anchor growing process. More implementation details are in Appendix C.4C.5.
4.4 Objective function
Watermark message loss.
As described in Sec.4.3, we use the pre-trained HiDDeN (Zhu et al., 2018) decoder to produce the watermark message $M^{\prime}$, and the watermark message loss $\mathcal{L}_{msg}$ is defined as a binary cross entropy (BCE) loss:
$\mathcal{L}_{msg}=\$$-\sum^{L}_{i=1}\Bigl[M_{i}\mathrm{log}(\sigma(M^{\prime}_{i}))+(1-M_{i})\mathrm{log}(1-\sigma(M^{\prime}_{i}))\Bigr],$ (6)
where $M\in\{0,1\}^{L}$ is the ground truth message, and a sigmoid function $\sigma$ constrains $M^{\prime}$ to $[0,1]$.
HSV Loss.
We incorporate the HSV loss $\mathcal{L}_{hsv}$ term to minimize color artifacts during watermark embedding. Previous studies (Huang et al., 2025; Jang et al., 2024; 2025) typically computed image loss in the RGB space to improve rendering quality. However, since the RGB space calculates loss based on independent pixel-level color information, it does not accurately reflect the human visual system. Therefore, we compute image loss in the HSV space, which represents colors in a manner more similar to the human visual system. First, we construct binary masks based on hue ranges for each color to localize color artifacts. Binary mask is defined as follows:
$M_{c}(p)=\begin{cases}1,&\text{if}\;\;\text{Hue}(p)\in H_{c}\\ 0,&\text{otherwise}\end{cases},$ (7)
where $H_{c}$ denotes the hue range corresponding to color $c\in\mathcal{C}$, with $\mathcal{C}=\{R,G,B\}$. $R$, $G$, and $B$ represent the red, green, and blue color channels, respectively. Then, HSV loss $\mathcal{L}_{\text{hsv}}$ is computed as the mean squared error (MSE) between the pixel value of rendered image $I(p)$ and the ground truth $I_{gt}(p)$ over the pixels in the set $\Omega$ where $M_{c}(p)=1$:
$\mathcal{L}_{\text{hsv}}=\frac{1}{|\mathcal{C}||\Omega|}\sum_{c\in\mathcal{C}}\sum_{p\in\Omega}\left\|M_{c}(p)\cdot\left(I(p)-I_{gt}(p)\right)\right\|^{2}.$ (8)
The overall color loss is defined as the average of the losses for each target color in the set $\mathcal{C}$. More implementation details are in Appendix C.6.
Figure 3: Qualitative comparison of rendering quality after compression between our method and baselines. For each method, the rendered image (left) and the difference map (right) are shown. Note that the difference maps are magnified five times. This result is based on 48-bit messages.
Total Loss.
Additionally, we employ frequency loss $\mathcal{L}_{freq}$ to reduce high-frequency distortions during watermark embedding. The overall training loss for our method is given by:
$\mathcal{L}_{total}=\lambda_{img}($ $\mathcal{L}_{scaffold}+\lambda_{hsv}\mathcal{L}_{hsv}+\lambda_{freq}\mathcal{L}_{freq})+\lambda_{msg}\mathcal{L}_{msg},$ (9)
where $\mathcal{L}_{freq}=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}P_{error}(p)$, as the mean of all pixel-wise SSIM losses. $\mathcal{L}_{scaffold}$ represents the reconstruction loss defined in Scaffold-GS (Lu et al., 2024), which consists of one scale regularization and two rendering fidelity terms.
5 Experiments
5.1 Experimental Setting
Datasets.
Following previous works Huang et al. (2025); Jang et al. (2025), we evaluate our method using the Blender Mildenhall et al. (2021), LLFF Mildenhall et al. (2019), and Mip-NeRF 360 Barron et al. (2022) datasets, comprising a total of 25 scenes. These include 8 synthetic bounded scenes from Blender, 8 forward-facing real-world scenes from LLFF, and 9 bounded real-world scenes from Mip-NeRF 360. We evaluate performance using 200 test images from Blender. For the LLFF and Mip-NeRF 360 datasets, we adopt the same data split strategy as Mip-NeRF 360.
Implementation details.
Our method trains end-to-end on a single A6000 GPU. We conduct experiments for watermark bit-lengths of 32, 48, and 64, focusing on the 48-bit case, where we evaluate performance both before and after compression. For the decoder, we use a pre-trained HiDDeN (Zhu et al., 2018) decoder for each bit-length and keep its parameters fixed during watermark training. We use the following parameters: $\lambda_{img}=10,\lambda_{hsv}=0.6,\lambda_{freq}=0.1,\lambda_{msg}=0.45$. For the quantization scale and anchor growing range in high-frequency regions, we set $\mathrm{Q}_{0}=1,\epsilon=0.3$. More implementation details are in Appendix C.3.
Table 1: Quantitative comparison of before (left) and after (right) compression bit accuracy and rendering quality. Evaluations are performed using a 48-bit setting, averaged over the Blender, LLFF, and Mip-NeRF 360 datasets. Baselines are tested within an anchor-based 3DGS framework with HAC and ContextGS compression. The best results are in bold.
Methods Bit Accuracy(%) ↑ PSNR ↑ SSIM ↑ LPIPS ↓ Size(MB) ↓ HAC + WateRF 91.02 / 54.40 27.36 / 13.63 0.850 / 0.457 0.174 / 0.574 207.72 / 13.66 HAC + GaussianMarker 92.00 / 58.34 27.05 / 13.54 0.840 / 0.460 0.193 / 0.571 341.30 / 24.33 HAC + 3D-GSW 90.96 / 53.48 19.57 / 13.13 0.628 / 0.439 0.295 / 0.572 173.64 / 12.93 HAC + GuardSplat 79.77 / 52.72 16.29 / 12.28 0.584 / 0.425 0.417 / 0.609 215.69 / 13.15 HAC + CompMarkGS 95.95 / 95.92 27.68 / 27.65 0.856 / 0.852 0.171 / 0.177 208.96 / 12.23 ContextGS + WateRF 92.01 / 90.36 26.64 / 26.47 0.843 / 0.832 0.183 / 0.185 219.48 /   9.88 ContextGS + GaussianMarker 91.24 / 87.95 26.89 / 26.54 0.839 / 0.827 0.195 / 0.200 342.38 / 17.86 ContextGS + 3D-GSW 88.28 / 79.75 19.50 / 19.83 0.627 / 0.617 0.294 / 0.299 177.19 /   9.24 ContextGS + GuardSplat 73.20 / 67.16 16.90 / 16.86 0.653 / 0.628 0.353 / 0.360 219.82 /   9.33 ContextGS + CompMarkGS 94.36 / 94.03 27.60 / 27.55 0.845 / 0.844 0.172 / 0.173 73.39 /   5.72
Figure 4: Robustness to image distortion. Based on 48-bit, the results are averaged over three datasets.
Table 2: Robustness to model distortion. Based on 48-bit, results are averaged over three datasets with both HAC&ContextGS before (left) and after (right) compression.
Bit Accuracy(%) $\uparrow$ Methods No Distortion Gaussian Noise ($\sigma$=0.005) Clone (50%) Prune (20%) WateRF 91.52 / 72.38 81.70 / 63.27 90.23 / 67.45 87.81 / 66.25 GaussianMarker 91.62 / 73.15 73.01 / 65.07 89.83 / 72.17 87.43 / 85.73 3D-GSW 89.62 / 62.89 78.56 / 57.72 88.26 / 62.14 84.84 / 59.76 GuardSplat 76.49 / 59.94 65.28 / 55.53 73.76 / 63.20 69.19 / 65.19 CompMarkGS 95.16 / 94.98 84.58 / 80.10 94.35 / 94.16 93.53 / 93.42
Baselines.
To validate our method, we evaluate watermark extraction before and after quantization-based compression in an anchor-based 3DGS framework. For comparison, we evaluate against the following four baselines: 1) WateRF (Jang et al., 2024): Embedding watermark into NeRF (Mildenhall et al., 2021) via frequency-domain (DWT) techniques; 2) GaussianMarker (Huang et al., 2025): Integrating watermark‐specific Gaussians into the 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) based on per-Gaussian uncertainty estimates; 3) 3D-GSW (Jang et al., 2025): Embedding the watermark message by selectively removing Gaussians with low impact on rendering quality; 4) GuardSplat (Chen et al., 2024b): Leveraging Spherical Harmonics (SH) coefficients to embed the message. All baselines originally designed for NeRF or 3DGS are adapted to the anchor-based 3DGS framework for our experiments. Notably, GuardSplat requires specific modifications due to architectural incompatibilities, which are detailed in Appendix D.1. We compare performance before and after compression using two quantization-based compression schemes: HAC (Chen et al., 2024a) and ContextGS (Wang et al., 2024a).
Evaluation.
We evaluate the watermark performance of our proposed method, CompMarkGS, and the baselines according to three aspects: 1) Fidelity: We measure PSNR, SSIM, and LPIPS (Zhang et al., 2018) to assess the rendering quality of watermarked models before and after compression by comparing rendered images to the original images. 2) Robustness: We evaluate bit accuracy under various distortions before and after compression. We evaluate performance under distortions for rendered images, including Gaussian noise, rotation, scaling, Gaussian blur, cropping, and JPEG compression. We assess robustness against Gaussian noise, cloning, and pruning attacks for watermarked models. 3) Capacity: We measure bit accuracy for 32, 48, and 64-bit message lengths.
5.2 Experimental results
Fidelity of before and after compression.
[…]
Robustness for the image distortion.
[…]
Figure 5: Qualitative comparison of rendered images with watermarks embedded into different anchor attributes. Rendered images are obtained with 48-bit watermark embedding.
Figure 6: Qualitative comparison of rendering quality with full method (ours), without FAG (left), without $\mathcal{L}_{hsv}$ (right). Rendered images are obtained with 48-bit watermark embedding.
Robustness for the model distortion.
[…]
Capacity.
Table 3: Comparison of bit accuracy for our method and baselines at 32, 48, and 64-bit. Results are averaged over three datasets with both HAC&\&ContextGS before compression.
Bit Accuracy(%) $\uparrow$ Methods 32 bits 48 bits 64 bits WateRF 92.46 91.52 88.21 GaussianMarker 95.07 91.62 79.81 3D-GSW 93.00 89.62 86.31 GuardSplat 79.64 76.49 72.83 CompMarkGS 96.52 95.16 91.29
Increasing a watermark message capacity leaves less embedding budget per bit and makes it more vulnerable to noise and compression, which lowers the bit accuracy. We test bit accuracy for message lengths of 32, 48, and 64 bits. As shown in Tab. 3, while the accuracy for all methods decreases with longer messages, our method maintains significantly higher bit accuracy. In stark contrast to other methods, its minimal performance drop even at a 64-bit messages demonstrates our method’s higher effective messages capacity and greater robustness.
Table 4: Ablations of watermark embedding into different anchor attributes. Results are reported for 48-bit messages and averaged over three datasets using HAC and ContextGS after compression.
HAC ContextGS Target Parameter Bit Acc (%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ Bit Acc (%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ Position 67.91 12.05 0.625 0.350 82.37 18.49 0.632 0.432 Scaling 65.06 19.35 0.615 0.323 64.68 13.40 0.466 0.474 Offsets 94.88 24.61 0.792 0.226 93.98 25.32 0.755 0.322 Anchor Feature (Ours) 95.92 27.65 0.852 0.177 94.03 27.55 0.844 0.173
Table 5: Ablation studies on frequency-aware anchor growing (FAG), quantization distortion layer (QDL), and HSV loss $\mathcal{L}_{hsv}$. Results are reported for 48-bit messages and averaged over three datasets under HAC and ContextGS after compression.
HAC ContextGS FAG QDL $\mathcal{L}_{hsv}$ Bit Acc(%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ Bit Acc(%)↑ PSNR ↑ SSIM ↑ LPIPS ↓ 90.67 26.44 0.827 0.196 87.63 25.56 0.812 0.223 $\checkmark$ $\checkmark$ 93.95 27.54 0.849 0.178 90.35 27.15 0.841 0.198 $\checkmark$ $\checkmark$ 90.75 26.75 0.844 0.182 88.09 25.04 0.801 0.229 $\checkmark$ $\checkmark$ 92.57 27.49 0.852 0.179 90.23 27.19 0.841 0.200 $\checkmark$ $\checkmark$ $\checkmark$ 95.92 27.65 0.852 0.177 94.03 27.55 0.844 0.173
5.3 Ablation Study
Anchor-based 3DGS watermark embedding.
We embed the watermark into different anchor attributes (position, scaling, and offsets) and compare the rendering quality and bit accuracy. As discussed in the previous section (Sec.4.1), since the watermark embedding feature shares the same dimensions as the anchor feature, when we embed the watermark into a different attribute, we set the watermark feature to match the dimensions of that specific attribute for the experiment. As shown in Fig.5 and Tab.4, embedding the watermark into anchor attributes such as position and scaling leads to significant quality degradation, as these attributes directly adjust the Gaussian pose and shape. In contrast, our method embeds the watermark into the anchor feature, seamlessly concealing it without significantly degrading scene information.
Effectiveness of our contributions.
We conduct an ablation study to analyze the individual contribution of our key components: frequency-aware anchor growing (FAG), quantization distortion layer (QDL), and the HSV loss $\mathcal{L}_{hsv}$. The experiments are performed under HAC and ContextGS compression settings, with the results reported in Tab.5. Removing QDL causes the most significant drop in bit accuracy, demonstrating its crucial role in preventing watermark loss during quantization. Removing FAG degrades both rendering quality and bit accuracy. This indicates that by growing anchors exclusively in high-frequency regions, FAG mitigates conflicts between low-frequency watermarks and quality enhancement algorithms, thereby securing both rendering quality and watermark robustness (See Fig.6). Finally, removing the HSV loss worsens the LPIPS score and introduces subtle color artifacts. This suggests that the HSV loss effectively mitigates color artifacts that are difficult to handle with an RGB-space loss function alone (See Fig.6). Effective watermarking must strike a balance among three competing goals: robustness, invisibility, and capacity. Our experiments show that FAG, QDL, and the HSV loss not only serve unique roles but also work complementarily to significantly improve this balance between bit accuracy and rendering quality in compressed environments.
6 Conclusion
We propose CompMarkGS, a compression‐robust watermarking method. By integrating a quantization distortion layer, we train the watermark to withstand quantization‐based compression, and we introduce a learnable watermark embedding feature for anchor‐based insertion, achieving both high security and high fidelity. Additionally, our frequency‐aware anchor growing and HSV loss preserve high‐quality rendering performance. To the best of our knowledge, CompMarkGS is the first compression‐robust watermarking framework for 3DGS, making it well-suited for protecting 3D assets on resource‐constrained devices.
Ethics Statement
Provenance.
We only used publicly released datasets from academic publications (Blender (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022)). Additionally, the Blender and LLFF datasets are distributed under the CC BY 3.0 Unported license, while the Mip-NeRF 360 dataset is available under the Apache License 2.0.
Anonymous.
The Blender (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022) datasets represent synthetic and real-world data, respectively. The Blender dataset consists of 3D object scenes, while LLFF and Mip-NeRF 360 comprise indoor and outdoor real-world scenes. None of these datasets contains any human subjects. The details of each dataset are as shown in Tab.6.
Table 6: Dataset composition used for evaluation. Blender is synthetic, whereas LLFF and Mip-NeRF 360 are real-world scenes, with no human subjects included.
Dataset Scene1 Scene2 Scene3 Scene4 Scene5 Scene6 Scene7 Scene8 Scene9 Blender chair drums ficus hotdog lego materials mic ship LLFF fern flower fortress horns leaves orchids room trex Mip-NeRF 360 bicycle bonsai counter flowers garden kitchen room stump treehill
The use of large language models
We used an LLM in a limited manner for grammar and style polishing only. It did not contribute to research ideation, experimental design, analysis. The authors take full responsibility for all content.
Appendix
Appendix A Broader impacts
Our proposed CompMarkGS is an invisible watermarking method for 3D assets, such as 3D Gaussian Splatting (3DGS), which remains recoverable even after aggressive compression. Our work has three key social impacts:
  • Security: CompMarkGS enables copyright protection of 3D assets distributed across diverse digital environments without compromising visual quality. This capability supports copyright protection and sustainable revenue generation across a wide range of 3D content creators.
  • Applicability: CompMarkGS ensures watermark robustness under high-ratio compression, enabling protected 3DGS to be reliably utilized on resource-constrained devices, such as mobile phones and head-mounted displays. These protected 3DGS can be safely deployed in virtual reality (VR), augmented reality (AR), medical imaging, and other 3D applications without compromising watermark integrity.
  • Sustainability: To the best of our knowledge, CompMarkGS is the first watermarking method to achieve watermark robustness against compression. Building on this work, it will provide a foundation for other researchers and practitioners to extend the approach, contributing to the advancement of watermarking research across a variety of compression formats.
Appendix B Limitations and future works.
B.1 Limitations
One limitation of our proposed method, CompMarkGS, is its reliance on a pre-trained decoder for the watermark extraction process. However, this dependency constitutes a manageable, one-time computational cost. The decoder is trained only once for a specific message length (e.g., 32, 48, or 64 bits) and can then be universally applied to any 3D Gaussian Splatting model without the need for model-specific fine-tuning. This design choice significantly enhances the method’s practicality and scalability, as the initial training effort is amortized over countless applications.
B.2 Future works
Looking ahead, we have identified several promising directions for future work. First, we plan to extend our method to operate independently of 3D Gaussian Splatting compression, which would broaden its applicability to a wider range of scenarios. Concurrently, we aim to develop a more lightweight watermarking architecture. This initiative will focus on reducing the computational overhead of both the embedding and decoding stages, making our solution even more suitable for real-time or resource-constrained environments.
Appendix C More implementation details
C.1 Target Compression Schemes for Robustness Evaluation
To test the robustness of our watermark, we use two prominent anchor-based 3DGS compression methods HAC (Chen et al., 2024a) and ContextGS (Wang et al., 2024a). Both techniques follow a similar pipeline, extracting features for each anchor, using MLPs to predict a context-based distribution, and then applying differentiable quantization and entropy coding. This approach significantly compresses the model size while minimizing any loss in rendering quality.
However, the information loss that occurs during quantization can damage or destroy a watermark. Therefore, our core objective is to achieve robustness specifically against this kind of quantization-based compression.
C.2 Overview of the CompMarkGS pipeline
In this section, we detail the core components of the CompMarkGS pipeline, covering the overall watermark training procedure.
Watermarking embedding stage.
The watermarking process begins by combining the original anchor features with a learnable watermark embedding feature to create watermarked anchor features. These features are then passed through a quantization distortion layer (QDL), which simulates distortions from quantization by injecting noise, resulting in the quantized watermarked anchor features. These final features are fed into the MLPs to predict Gaussian parameters and render the scene. To recover the watermark message from the rendered image, we apply a Discrete Wavelet Transform (DWT) to extract the low-frequency band and input it into a pre-trained decoder. The entire training process is optimized to minimize the message reconstruction loss $\mathcal{L}_{msg}$, with an additional HSV-based perceptual loss $\mathcal{L}_{hsv}$ used to prevent visual quality degradation.
Frequency-aware anchor growing stage.
To further enhance rendering quality, we introduce the frequency-aware anchor growing (FAG) stage, which leverages frequency information to grow new anchors. We apply a Discrete Fourier Transform (DFT) and a high-pass filter to both the rendered and ground truth images to isolate their high-frequency components. A pixel-wise SSIM error map is then calculated between these two high-frequency representations. Based on this error map, we identify Gaussians in regions where high-frequency information is poorly reconstructed and use backpropagation to selectively grow new anchors in those locations.
In summary, CompMarkGS learns watermark robustness against compression-induced information loss via the QDL, while simultaneously preserving high-frequency details that could be degraded by watermarking via FAG. As a result, our proposed method achieves high watermark recovery rates and excellent visual fidelity in scenarios both with and without model compression.
C.3 Details of training
Our proposed CompMarkGS trains the anchor-based 3DGS from scratch with the watermark, in contrast to existing methods Huang et al. (2025); Jang et al. (2024; 2025) that rely on fine-tuning pre-trained 3DGS models. We train CompMarkGS for $30,000$ iterations. Anchor growing is performed between $1,600$ and $15,000$ iterations, first using conventional anchor growing and then switching to the proposed frequency-aware anchor growing strategy for fine-grained refinement. The loss function for watermark training consists of the original anchor-based 3DGS Lu et al. (2024) loss $\mathcal{L}_{\text{scaffold}}$ and three additional components: HSV loss $\mathcal{L}_{\text{hsv}}$, frequency loss $\mathcal{L}_{\text{freq}}$, and message loss $\mathcal{L}_{\text{msg}}$. Before watermark training begins, we pre-train a HiDDeN Zhu et al. (2018) message decoder on the MS-COCO dataset Lin et al. (2014) for 32, 48, and 64-bit messages following the strategy from 3D-GSW Jang et al. (2025) and Stable Signature Fernandez et al. (2023). During the subsequent watermark training, the pre-trained decoder is kept fixed. Additionally, the compression ratio of the compression method used in all experiments is set to $0.004$.
C.4 Details of high-pass filter
Early digital watermarking studies hide the watermark in the high-frequency bands of an image to keep it imperceptible. However, the watermark embedded in high-frequency regions is vulnerable to image distortions such as JPEG compression. More recent work, therefore, embeds the watermark in the low-frequency bands, which are more robust to such attacks. The drawback is that low-frequency components represent essential information, including the global structure and color distribution of the image. Modifying these components can degrade visual coherence and affect the perception of fine details such as edges and textures.
The conventional anchor growing strategy computes the mean gradient of the Gaussians within each voxel and places a new anchor point only if no anchor already exists in voxels whose mean gradient exceeds a threshold. During anchor-based 3DGS training, voxels with high mean gradients are primarily associated with regions that form the coarse structure of the scene. As a result, densification rarely occurs in fine-detail areas, and when a watermark is embedded, those under-represented regions are more susceptible to quality degradation.
To address this limitation, we propose a frequency-aware anchor growing strategy that selectively densifies Gaussians located in high-frequency regions. Specifically, we apply a Discrete Fourier Transform (DFT) to convert the rendered image into the frequency domain. To isolate high-frequency components, we use a high-pass filter mask $M_{h}$ that emphasizes high-frequency regions while suppressing low-frequency ones in the frequency domain:
$M_{h}(p)=1-\exp\biggl(-\frac{(d(p)-\tau)^{2}}{2\beta}\biggl),$ (10)
where $d(p)$ denotes the Euclidean distance between a pixel $p$ and the center of the image. $\tau$ defines the cutoff threshold of the mask and $\beta$ determines the degree of attenuation. The resulting mask $M_{h}$ is multiplied element-wise with the Fourier-transformed image to suppress low-frequency components and preserve only the high-frequency details. Afterward, we apply the Inverse Discrete Fourier Transform (IDFT) to reconstruct the filtered images in the spatial domain, resulting in the high-frequency rendered images $I^{\prime}_{hf}$ and the high-frequency ground truth images $I_{hf}$. Further details of the frequency-aware anchor growing process can be found in the main paper (See main paper Sec.4.3).
C.5 Details of frequency-aware anchor growing
We propose a frequency-aware anchor growing strategy that efficiently identifies and densifies Gaussians in high-frequency regions to mitigate visual quality degradation caused by watermark embedding. Specifically, we first compute a pixel-wise SSIM-based error map $P_{error}$ (See main paper Sec.4.3) between the high-frequency rendered image $I^{\prime}_{hf}$ and the ground truth image $I_{hf}$, both extracted in the frequency domain (See Sec.C.4). This error map $P_{error}$ highlights pixel regions where high-frequency information is not faithfully reconstructed. To identify pixel regions that contain visually important fine details, we compute the median value $\tilde{P}_{error}$ of the SSIM-based pixel-wise error map $P_{error}$. Using the median value $\tilde{P}_{error}$ ensures a more meaningful identification of high-frequency regions by mitigating the influence of outlier pixels, which may otherwise dominate when relying on the maximum value of the SSIM-based pixel-wise error map $P_{error}$. We then generate a binary mask $I_{mask}$ by filtering pixels whose error values fall within the interval $\tilde{P}_{error}-\epsilon,\tilde{P}_{error}+\epsilon]$:
$I_{mask}(p)=\begin{cases}1,&\text{if}\;\;|P_{error}(p)-\tilde{P}_{error}|\leq\epsilon\\ 0,&\text{otherwise}\end{cases},$ (11)
where $p$ denotes the pixel coordinates in the image. Subsequently, we construct a binary mask $G_{mask}$ to select the 3D Gaussian points that 2D projections fall within the boundaries of the image:
$G_{mask}(k)=\begin{cases}1,&\text{if}\;\;0\leq x<W\quad\mathrm{and}\quad 0\leq y<H\\ 0,&\text{otherwise}\end{cases},$ (12)
where $k=(x,y)$ denotes the pixel coordinates of the 2D Gaussian, and $W$ and $H$ represent the width and height of the image, respectively. Using the two binary masks $I_{mask}$ and $G_{mask}$, we generate a boolean mask $F_{mask}$ that identifies Gaussians located in specific high-frequency regions by matching their 2D coordinates to the high-frequency pixel locations:
$F_{mask}(k)=I_{mask}(k)\cdot G_{mask}(k),$ (13)
where $k$ denotes the pixel coordinates of the 2D Gaussian. During the anchor growing process, we calculate the average of Gaussians within the voxel regions selected by the computed boolean mask $F_{mask}$, based on the accumulated sum of their gradients. If the average gradient exceeds the threshold $\tau^{g}_{high}=0.00015$ and no anchor exists in the corresponding voxel, a new anchor is added. The threshold for the accumulated opacity used in the subsequent anchor pruning step is set to $\tau^{o}_{high}=0.15$.
C.6 Details of HSV loss
In the field of digital watermarking, achieving imperceptible watermark embedding is a primary research goal. However, there exists an inherent trade-off between invisibility and bit accuracy. When training a model to embed a watermark in 3D Gaussian Splatting (3DGS), color artifacts emerge due to the tendency to embed the watermark into specific Gaussians to achieve high bit accuracy. Additionally, these artifacts occur particularly in regions rendered by a small number of Gaussians, where even slight modifications to the Gaussians can lead to noticeable distortions.
Although the RGB color space is widely used for image representation, prior work on image sharpening Kau & Lee (2013) indicates that the RGB color space does not fully account for the perceptual characteristics of the human visual system. In particular, the human visual system is more sensitive to luminance variations than to chromatic changes Gonzalez (2009). Unlike the RGB color space, the Hue, Saturation, and Value (HSV) space separates chromatic information (Hue and Saturation) from luminance (Value), enabling more precise detection and suppression of color artifacts. Based on this property, we introduce an HSV loss that enhances rendering quality while preserving high bit accuracy.
To isolate specific regions in the HSV space, we construct binary masks (See main paper Eq.7) by filtering pixels based on predefined hue ranges $H_{c}$, along with threshold conditions on saturation $\tau_{c}^{s}$ and value $\tau_{c}^{v}$, corresponding to color $c\in\mathcal{C}$, where $\mathcal{C}=\{R,G,B\}$. Each pixel $p\in\Omega$, where $\Omega\subset\mathbb{R}^{2}$ denotes the spatial domain of the image, is evaluated based on its HSV values. Let $S(p)\in[0,1]$ and $V(p)\in[0,1]$ denote the saturation and value components at pixel $p$, respectively. We define the hue range $H_{c}$ for each color as follows:
$H_{R}\in\left[0,\dfrac{\pi}{3}\right)\cup\left[\dfrac{5\pi}{3},2\pi\right),\;\;\text{if}\;\;S(p)\geq\tau_{R}^{s}\;\;\text{and}\;\;V(p)\geq\tau_{R}^{v},$ (14)
$H_{G}\in\left[\dfrac{\pi}{3},\pi\right),\;\;\text{if}\;\;S(p)\geq\tau_{G}^{s}\;\;\text{and}\;\;V(p)\geq\tau_{G}^{v},$
$H_{B}\in\left[\pi,\dfrac{5\pi}{3}\right),\;\;\text{if}\;\;S(p)\geq\tau_{B}^{s}\;\;\text{and}\;\;V(p)\geq\tau_{B}^{v},$
where $H(p)\in[0,2\pi)$ denotes the hue value at pixel $p$. Pixels that do not satisfy the corresponding saturation or value thresholds are excluded from the mask, even if their hue lies within the specified range. We set thresholds for the red with $\tau_{R}^{s}=0.4$ and $\tau_{R}^{v}=0$, while using relaxed thresholds $\tau_{c}^{s}=0.2$, $\tau_{c}^{v}=0.2$ for the green and blue.
Appendix D Discussion
D.1 Comparison with related work
Distinction from anchor-based approaches.
Although both CompMarkGS and SecureGS use an anchor-based structure, they are fundamentally different, starting from their problem definition. SecureGS (Zhang et al., 2025b) is a steganography method aiming to hide large amounts of data, with invisibility as its key metric. In contrast, CompMarkGS is a digital watermarking method for proving ownership, which prioritizes robustness. Notably, the ”robustness” evaluation in SecureGS focuses on preserving the rendering quality of the hidden scene, whereas CompMarkGS aims for the recovery of the hidden message itself under attacks like model compression. To achieve these different goals, their information embedding mechanisms are also distinct. SecureGS predicts hidden Gaussian offsets with a private MLP, while CompMarkGS directly injects a learnable watermark embedding feature into the anchor feature and extracts the message from the rendered image via a pre-trained decoder. Therefore, CompMarkGS is not an incremental improvement on SecureGS, but a unique solution independently designed for the new objective of verifiable ownership after model compression.
Comparison with 3DGS watermarking techniques.
A key difference between our method and existing watermarking methods, particularly frequency-based approaches such as 3D-GSW (Jang et al., 2025), lies in our unique approach to utilizing frequency information. While both methods leverage frequency data for Gaussian densification, their core strategies and objectives are fundamentally different. 3D-GSW measures the frequency intensity of patches, whereas CompMarkGS precisely identifies high-frequency error regions between the rendered image and the ground truth, selectively densifying anchors in those specific areas. However, the most significant distinction is the role of compression in the watermarking pipeline. Unlike 3D-GSW, which incorporates model pruning as part of its watermarking process, CompMarkGS is the first to define and solve the novel problem of watermark robustness against external quantization compression applied to a fully trained model. To achieve this, we introduce a quantization distortion layer (QDL) that simulates quantization noise during training, a core component absent in 3D-GSW. Our contribution is proposing a unique method specifically optimized to address the previously unexplored challenge of watermark loss during quantization compression. This focus on compression robustness also sets CompMarkGS apart from GaussianMarker (Huang et al., 2025) and GuardSplat (Chen et al., 2024b). These methods address traditional image distortions like Gaussian noise and JPEG compression, or model distortions such as pruning. However, they lack a mechanism to mitigate the loss of watermark information caused by quantization compression.
Furthermore, for a fair comparison, we adapt the existing vanilla 3DGS-based watermarking techniques to the anchor-based 3DGS framework. However, during this process, GuardSplat requires significant adaptation due to a core architectural incompatibility. This is because GuardSplat relies on adding offsets to explicit Spherical Harmonics (SH) parameters, which do not exist in our anchor-based model. Therefore, we modified its mechanism to apply offsets to the anchor features instead for our experiments.
D.2 Motivation for watermarking anchor-based 3D Gaussian Splatting
As the demand for lightweight and efficiently rendered 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) models grows, anchor-based architectures are emerging as a key solution. In particular, models like Scaffold-GS (Lu et al., 2024), which combine anchors with a MLP, achieve both high memory efficiency and excellent rendering quality compared to vanilla 3DGS. This anchor-based structure is highly practical, making it ideal for mobile and AR/VR environments where real-time rendering is essential.
We identified these anchor-based 3DGS models as ideal targets for watermarking due to a significant security advantage. Instead of directly storing and rendering Gaussian attributes, anchor-based models generate them indirectly by feeding anchor features into an MLP. If we embed a watermark into these anchor features, it undergoes a non-linear transformation by the MLP. This process expresses the watermark within the Gaussian attributes in a complex and concealed manner, making it extremely difficult for an external attacker to directly interpret or remove it.
Therefore, selecting an anchor-based 3DGS model for watermarking is a rational approach that secures both practical benefits, such as a lightweight structure, and robust security through the MLP.
Appendix E Additional experimental results
E.1 Effectiveness of watermarking embedding feature
To verify the effectiveness of our proposed watermark embedding strategy, we compare the performance of two approaches: using only the anchor feature and using a dedicated watermark embedding feature. As described in Sec.4.1, our method introduces a dedicated watermark embedding feature that shares the same dimension as the anchor feature, combining them via element-wise sum. This design ensures that the total feature dimension remains constant, regardless of whether a watermark is embedded. Therefore, any observed performance increase can be attributed to the effectiveness of the embedding strategy itself, rather than an increase in model capacity.
Table 7: Ablation study on the dedicated watermark embedding feature. Results are reported for 48-bit messages and averaged over three datasets after ContextGS compression.
Embedding Method Bit Acc. (%) ↑ PSNR ↑ SSIM ↑ LPIPS ↓ Anchor Feature (Ours w/o $f^{\prime}$) 92.94 26.79 0.834 0.209 Anchor Feature (Ours w/ $f^{\prime}$) 94.03 27.55 0.844 0.173
As shown in Tab.7, using only the anchor feature creates a conflict between the two tasks of representing the scene and hiding information, which leads to degraded rendering quality. In contrast, our approach of introducing a dedicated watermark embedding feature preserves high rendering quality while enabling a more robust watermark. We conclude that using an independent watermark embedding feature is essential for effectively decoupling these two objectives, thereby achieving an optimal balance between rendering quality and watermark robustness.
E.2 Effectiveness of HSV loss
When only RGB-based loss functions are used for watermark embedding, the optimization of 3DGS fails to consider the characteristics of the human visual system. As a result, 3DGS with a watermark renders color artifacts in the rendered images, particularly in regions with strong chromatic and luminance variations. These artifacts are observed in Fig.7 and Fig.8, despite achieving high bit accuracy. To address this, we introduce an HSV loss that aligns more closely with the perceptual characteristics of the human visual system. HSV color space separates chromatic components (Hue and Saturation) from luminance (Value), making it more suitable for detecting and reducing color artifacts. As shown in Fig. 7, binary masks of HSV loss successfully detect color artifacts in each color channel. Moreover, Fig. 8 shows that HSV loss allows 3DGS to focus its optimization on minimizing color artifacts, maintaining high bit accuracy. These results show that HSV loss effectively reduces color artifacts while preserving high bit accuracy and rendering quality.
Figure 7: Visualization of the binary mask and masked image for each color channel (Red, Green, and Blue). The left column shows the ground truth and the rendered image without HSV loss. The right columns present the binary masks (top) and masked results (bottom) for each color channel. This result is based on 48-bit messages.
Figure 8: Visualization of HSV loss. The comparison is performed between images rendered with and without the HSV loss. This result is based on 48-bit messages.
E.3 Robustness under various conditions
Robustness to pruning-based compression.
[…]
Method Bit Acc. (%) ↑ Size (MB) ↓ WateRF 92.65 / 67.29 248.97 / 84.77 GaussianMarker 92.89 / 66.67 418.66 / 141.84 3D-GSW 91.93 / 67.06 228.62 / 77.46 CompMarkGS 96.09 / 83.50 186.71 / 61.61
Robustness to general quantization.
[…]
Table 9: Robustness comparison against general quantization, showing the bit accuracy drop on the Mip-NeRF 360 dataset (48-bit messages) after quantizing models from float32 to int8.
Method Bit Acc. (% )↑ PSNR ↑ SSIM ↑ LPIPS ↓ WateRF 93.20 / 69.20 26.09 / 12.17 0.788 / 0.135 0.260 / 0.610 GaussianMarker 92.49 / 66.02 26.33 / 15.83 0.775 / 0.407 0.282 / 0.588 3D-GSW 91.08 / 66.20 18.53 / 14.93 0.667 / 0.374 0.322 / 0.662 CompMarkGS 95.59 / 90.11 26.56 / 21.73 0.777 / 0.700 0.248 / 0.336
E.4 Generality and Compatibility
Robustness to adversarial attacks.
[…]
Table 10: Robustness comparison against PGD adversarial attacks. Results are for 48-bit messages, showing bit accuracy before and after the attack targeting the HiDDeN decoder.
Method Bit Acc. (None)↑ Bit Acc. (PGD $\epsilon=0.1$) ↑ WateRF 92.01 56.06 GaussianMarker 87.95 57.53 3D-GSW 79.25 57.56 CompMarkGS 94.03 84.33
Decoder independence.
[…]
Table 11: Performance comparison with an alternative decoder (DINO-based SSL). Results are for 48-bit messages, averaged over three datasets, showing performance before and after HAC and ContextGS compression.
Methods Bit Accuracy (%) ↑ PSNR ↑ SSIM ↑ LPIPS ↓ HAC + WateRF 91.02 / 54.40 27.36 / 13.63 0.850 / 0.457 0.174 / 0.574 HAC + GaussianMarker 92.00 / 58.34 27.05 / 13.54 0.840 / 0.460 0.193 / 0.571 HAC + 3D-GSW 90.96 / 53.48 19.57 / 13.13 0.628 / 0.439 0.295 / 0.572 HAC + GuardSplat 79.77 / 52.72 16.29 / 12.28 0.584 / 0.425 0.417 / 0.609 HAC + CompMarkGS (SSL) 93.54 / 93.50 27.21 / 27.16 0.847 / 0.842 0.190 / 0.192 HAC + CompMarkGS (HiDDeN) 95.95 / 95.92 27.68 / 27.65 0.856 / 0.852 0.171 / 0.177 ContextGS + WateRF 92.01 / 90.36 26.64 / 26.47 0.843 / 0.832 0.183 / 0.185 ContextGS + GaussianMarker 91.24 / 87.95 26.89 / 26.54 0.839 / 0.827 0.195 / 0.200 ContextGS + 3D-GSW 88.28 / 79.75 19.50 / 19.83 0.627 / 0.617 0.294 / 0.299 ContextGS + GuardSplat 73.20 / 67.16 16.90 / 16.86 0.653 / 0.628 0.353 / 0.360 ContextGS + CompMarkGS (SSL) 93.09 / 93.00 25.86 / 25.82 0.829 / 0.828 0.220 / 0.221 ContextGS + CompMarkGS (HiDDeN) 94.36 / 94.03 27.60 / 27.55 0.845 / 0.844 0.172 / 0.173
E.5 Robustness to image distortion comparison
This section provides the detailed quantitative results for the experiment on robustness to image distortion, which is visually summarized in Fig.4 of the main paper. We evaluate watermark resilience by applying six post-processing distortions to the rendered images: Gaussian noise, rotation, scaling, Gaussian blur, crop, and JPEG compression. Tab.12 details the bit accuracy for each distortion condition, both before (left) and after (right) compression. As the results show, our proposed method, CompMarkGS, consistently exhibits superior robustness to existing methods across most distortion and compression scenarios.
Table 12: Robustness to image distortion before (left) and after (right) compression. Evaluations are performed using a 48-bit setting, averaged over the Blender, LLFF, and Mip-NeRF 360 datasets. Baselines are tested within an anchor-based 3DGS framework with HAC and ContextGS compression. The best results are in bold.
Bit Accuracy(%) $\uparrow$ Methods No Distortion Gaussian Noise ($\sigma$ = 0.1) Rotation ($\pm\pi/6$) Scaling (75%) Gaussian Blur ($\sigma$ = 0.1) Crop (40%) JPEG Compression (50% quality) HAC + WateRF 91.02 / 54.40 89.75 / 55.90 84.92 / 56.16 86.17 / 55.38 89.33 / 56.34 87.00 / 55.30 84.00 / 57.47 HAC + GaussianMarker 92.00 / 58.34 59.08 / 58.33 58.75 / 58.33 58.33 / 58.33 59.08 / 58.33 61.42 / 58.17 58.25 / 58.33 HAC + 3D-GSW 90.96 / 53.48 89.50 / 52.50 86.67 / 52.25 86.42 / 52.08 86.67 / 52.25 86.92 / 52.50 83.83 / 53.25 HAC + GuardSplat 79.77 / 52.72 80.67 / 50.17 74.42 / 53.33 81.33 / 50.75 81.50 / 50.33 65.08 / 50.00 74.83 / 52.67 HAC + CompMarkGS 95.95 / 95.92 96.00 / 95.50 92.25 / 92.08 91.83 / 91.67 95.83 / 95.58 90.67 / 90.83 87.75 / 88.83 ContextGS + WateRF 92.01 / 90.36 92.25 / 91.50 88.75 / 88.00 89.58 / 88.83 92.75 / 91.50 89.50 / 88.25 86.41 / 84.58 ContextGS  + GaussianMarker 91.24 / 87.95 59.08 / 58.92 59.00 / 58.75 58.25 / 58.17 59.17 / 58.92 61.08 / 60.67 58.25 / 58.25 ContextGS  + 3D-GSW 88.28 / 79.25 87.92 / 80.67 85.08 / 78.75 84.42 / 77.08 88.00 / 81.42 84.33 / 78.50 81.58 / 72.58 ContextGS  + GuardSplat 73.20 / 67.16 73.92 / 65.83 67.00 / 61.83 74.25 / 66.67 74.33 / 67.08 62.33 / 54.75 63.75 / 61.67 ContextGS + CompMarkGS 94.36 / 94.03 93.04 / 92.98 90.17 / 90.33 90.46 / 90.06 92.96 / 92.81 89.63 / 89.29 85.21 / 85.31
E.6 Robustness to model distortion comparison
Fig.9(a) and Fig.9(b) show the bit accuracy under different levels of model distortion. We conduct three model distortions: 1) adding Gaussian noise to all model parameters, 2) randomly removing anchors, and 3) randomly cloning anchors. Across different distortion strengths, our proposed method consistently outperforms the baseline. Notably, under prune and clone distortion settings, our method exhibits less performance degradation, owing to the incorporation of both pruning and anchor growing during training in CompMarkGS. These results demonstrate that our method achieves superior performance over the baseline, even under severe model distortions.
(a) HAC compression.
(b) ContextGS compression.
Figure 9: Comparison of baselines and our method under Gaussian noise, cloning, and pruning. Results are averaged over Blender, LLFF, and Mip-NeRF 360 with 48-bit messages.
Table 13: Robustness to model distortion before (left) and after (right) compression.
Bit Accuracy(%) $\uparrow$ Methods No Distortion Gaussian Noise ($\sigma$ = 0.005) Clone (50%) Prune (20%) HAC + WateRF 91.02 / 54.40 81.03 / 53.38 89.78 / 54.04 87.33 / 53.70 HAC + GaussianMarker 92.00 / 58.34 73.18 / 58.35 90.23 / 58.29 87.90 / 87.89 HAC + 3D-GSW 90.96 / 53.48 80.17 / 53.11 89.62 / 53.17 86.80 / 53.18 HAC + GuardSplat 79.77 / 52.72 67.32 / 51.21 76.12 / 56.28 71.25 / 59.13 HAC + CompMarkGS 95.95 / 95.92 82.13 / 77.13 95.33 / 95.09 94.67 / 94.25 ContextGS + WateRF 92.01 / 90.36 82.36 / 79.91 90.70 / 88.95 88.29 / 86.20 ContextGS  + GaussianMarker 91.24 / 87.95 72.84 / 71.78 89.43 / 86.05 86.97 / 83.56 ContextGS  + 3D-GSW 88.28 / 79.75 76.94 / 68.55 86.89 / 77.99 82.87 / 72.70 ContextGS  + GuardSplat 73.20 / 67.16 63.24 / 59.84 71.39 / 70.11 67.12 / 65.19 ContextGS + CompMarkGS 94.36 / 94.03 87.03 / 83.08 93.37 / 93.23 92.38 / 92.57
E.7 Performance across compression levels
Compression plays a crucial role in determining the trade-off between model size and performance in terms of bit accuracy and rendering quality. Fig.10 shows the relationship among model size, bit accuracy, and rendering quality under two different compression schemes: HAC (Chen et al., 2024a) and ContextGS (Wang et al., 2024a). Fig.10 indicates that a lower compression level leads to better bit accuracy and rendering quality. Our method consistently outperforms other methods, achieving the highest bit accuracy and rendering quality across varying model sizes. Notably, for HAC compression, we achieve bit accuracy of $90\%$ with model size less than $10$ MB, while other methods only reach $60\%$ under the same model size. These results highlight the effectiveness and robustness of CompMarkGS.
(a) HAC compression.
(b) ContextGS compression.
Figure 10: Performance of bit accuracy and rendering quality under different compression levels. A larger model size corresponds to a lower compression level. The blue line represents the results of our method. Results represent the average score across Blender, LLFF, and Mip-NeRF 360 datasets using 48-bit messages.
Appendix F Computing resources
We evaluate the rendering performance of our method and baselines (Huang et al., 2025; Jang et al., 2024; 2025) by measuring FPS. As shown in Tab.14, our method exceeds the real-time rendering threshold of $30$ FPS across the Blender (Mildenhall et al., 2021), LLFF (Mildenhall et al., 2019), and Mip-NeRF 360 (Barron et al., 2022) datasets. These results demonstrate the practical applicability and rendering efficiency of our proposed CompMarkGS.
Table 14: FPS results for each of the Blender, LLFF, and Mip-NeRF 360 datasets. Based on 48-bit messages, results are averaged with both HAC and ContextGS after compression.
FPS ↑ Methods Blender LLFF Mip-NeRF 360 WateRF 199.30 19.31 41.98 GaussianMarker 166.45 16.24 28.54 3D-GSW 220.60 22.67 50.64 GuardSplat 207.56 21.93 41.28 CompMarkGS 230.39 35.30 67.58
Appendix G Additional qualitative results
Fig.11 through Fig.19 visualize all results rendered after compression using our method with HAC and ContextGS, along with the difference (×5) between the original images and the watermarked images.
Figure 11: Rendering quality of various rendering outputs generated by our method on the Mip-NeRF 360 dataset. We show the differences (× 5). The closer it is to white, the greater the discrepancy between the ground truth and the rendered image. The results were obtained using 32-bit messages after compression.
Figure 12: Rendering quality of various rendering outputs generated by our method on the LLFF dataset. We show the differences (× 5). The closer it is to white, the greater the discrepancy between the ground truth and the rendered image. The results were obtained using 32-bit messages after compression.
Figure 13: Rendering quality of various rendering outputs generated by our method on the Blender dataset. We show the differences (× 5). The closer it is to white, the greater the discrepancy between the ground truth and the rendered image. The results were obtained using 32-bit messages after compression.
Figure 14: Rendering quality of various rendering outputs generated by our method on the Mip-NeRF 360 dataset. We show the differences (× 5). The closer it is to white, the greater the discrepancy between the ground truth and the rendered image. The results were obtained using 48-bit messages after compression.
Figure 15: Rendering quality of various rendering outputs generated by our method on the LLFF dataset. We show the differences (× 5). The closer it is to white, the greater the discrepancy between the ground truth and the rendered image. The results were obtained using 48-bit messages after compression.
Figure 16: Rendering quality of various rendering outputs generated by our method on the Blender dataset. We show the differences (× 5). The closer it is to white, the greater the discrepancy between the ground truth and the rendered image. The results were obtained using 48-bit messages after compression.
Figure 17: Rendering quality of various rendering outputs generated by our method on the Mip-NeRF 360 dataset. We show the differences (× 5). The closer it is to white, the greater the discrepancy between the ground truth and the rendered image. The results were obtained using 64-bit messages after compression.
Figure 18: Rendering quality of various rendering outputs generated by our method on the LLFF dataset. We show the differences (× 5). The closer it is to white, the greater the discrepancy between the ground truth and the rendered image. The results were obtained using 64-bit messages after compression.
Figure 19: Rendering quality of various rendering outputs generated by our method on the Blender dataset. We show the differences (× 5). The closer it is to white, the greater the discrepancy between the ground truth and the rendered image. The results were obtained using 64-bit messages after compression.